### v1

# Surface charge density wave in UTe2

Surface charge density wave in UTe2
Pablo García Talavera Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Miguel Águeda Velasco Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Makoto Shimizu Department of Physics, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan. Beilun Wu Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Georg Knebel University Grenoble Alpes, CEA, IRIG-PHELIQS, F-38000 Grenoble, France. Midori Amano Patino University Grenoble Alpes, CEA, IRIG-PHELIQS, F-38000 Grenoble, France. Univ. Grenoble Alpes, CNRS, Grenoble INP, Institut Néel, 38000 Grenoble, France Univ. Grenoble Alpes, CEA, IRIG, Pheliqs, 38000 Grenoble, France. Gerard Lapertot University Grenoble Alpes, CEA, IRIG-PHELIQS, F-38000 Grenoble, France. Jacques Flouquet Institute for Materials Research, Tohoku University, Ibaraki 311-1313, Japan. Jean Pascal Brison University Grenoble Alpes, CEA, IRIG-PHELIQS, F-38000 Grenoble, France. Dai Aoki Institute for Materials Research, Tohoku University, Ibaraki 311-1313, Japan. Youichi Yanase University Grenoble Alpes, CEA, IRIG-PHELIQS, F-38000 Grenoble, France. Edwin Herrera Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Isabel Guillamón Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Hermann Suderow Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain.
The spatially uniform electronic density characteristic of a metal can become unstable at low temperatures, leading to the formation of charge density waves (CDWs). These CDWs, observed in dichalcogenides, cuprates, and pnictides arise from features in the electron and lattice bandstructures that facilitate charge ordering. CDWs are often considered to compete with Kondo screening and are relatively rare in heavy fermion metals. However, the heavy fermion topological superconductor candidate UTe2 presents a notable exception, exhibiting a CDW whose origin remains elusive. Here we report high resolution Scanning Tunneling Microscopy (STM) experiments that reveal the primitive wavevectors of the CDW in UTe2. This allows for a refined identification of the electronic bandstructure regions susceptible to nesting. We demonstrate that the CDW wavevectors are not linked to bulk antiferromagnetic fluctuations that have been connected to other nesting features, indicating a decoupling from the bulk. We propose that surface-induced modifications of the U-5f electronic structure result in an enhancement of electronic interactions specifically at the nesting wavevectors identified here, thereby driving the formation of the observed surface CDW.
The CDW in UTe2.
The unconventional superconductivity observed in UTe2 exhibits characteristics strongly suggestive of spin-triplet Cooper pairing. The superconducting phase diagram, featuring an exceptionally high critical field and reentrant behavior at extreme magnetic fields [1, 2, 3], anomalous field dependence of the Knight-shift [4, 5] and multiple transitions between superconducting phases [6, 7, 8, 9, 10, 11, 12, 4], poses significant challenges to explanations based on conventional s-wave BCS theory. A relevant question concerns the nature of the normal state parent to the remarkable superconducting properties of UTe2.
Scanning Tunneling Microscopy (STM) studies have revealed the presence of a charge density wave (CDW) [13, 14, 15, 16, 17, 18]. Analogous to the cuprates [19, 20, 21], the CDW is linked to a spatially modulated superconducting phase, termed pair density wave (PDW) [15, 22]. However, the underlying mechanism linking the CDW to electronic and lattice properties remains poorly understood. Notably, structural studies have shown the absence of detectable lattice distortions characteristic of conventional CDW formation [23, 24], and macroscopic measurements, such as specific heat or thermal expansion have failed to identify signatures of a phase transition associated with the CDW [25, 26]. Moreover, the observed CDW wavevectors have apparently no relationship with features of the bulk band structure [14, 15, 16, 17, 27, 28, 29, 30, 31, 32, 33]. The fact that the hitherto observed CDW wavevectors lie outside the first Brillouin zone further complicates the identification of its connection to bulk electronic properties. Here, we report new high resolution millikelvin and high magnetic field STM measurements demonstrating that the previously identified wavevectors are not primitive wavevectors of the CDW lattice.
Primitive wavevectors of the CDW of UTe2.
The cleaved surface of UTe2, depicted in grey in Fig. 1a, corresponds to the (011) plane of the orthorhombic crystal structure (space group #71, $Immm$). The STM topography image in Fig. 1b reveals chains formed by Te(2) atoms, with chains of Te(1) located slightly below. Uranium (U) chains are located between the Te(1) and Te(2) surface chains. The atomic arrangement at the surface defines the surface centered rectangular unit cell, indicated by the solid blue lines in Fig. 1b. The primitive crystal surface unit cell is represented by dashed blue lines in Fig. 1b. We applied a magnetic field perpendicular to the surface, along a direction within the superconducting phase diagram at $23.7^{\circ}$ from the b axis towards the c axis direction, as shown in Fig. 1c. This magnetic field direction aligns with the onset of the very high field superconducting phase above 35 T [1, 2, 3]. The bulk Brillouin zone, with standard high symmetry point labels, is shown in black lines in Fig. 1d-f. The previously reported CDW Bragg peaks are shown in orange, and lie outside of both the bulk and the surface first Brillouin zones (detailed in Extended Data Fig. 1 and Supplementary Information, Section 1). In contrast, the Bragg peaks observed in this study, depicted in yellow, are located within the first Brillouin zones of both the bulk and surface lattices.
We present atomic resolution tunneling conductance maps acquired near zero bias from 10 T to 20 T in Fig. 2a-c, illustrating data from three different fields of view. The corresponding Fourier transforms are shown in Fig. 2d-f. Tunneling conductance maps at different bias voltages are provided in Extended Data Fig. 2 and in Supplementary videos. We identify Bragg peaks associated with the atomic lattice and an additional set corresponding to the CDW. The surface Brillouin zone is delineated by a red hexagon in Fig. 2d-f. The CDW peaks indicated by orange circles are consistent with those reported in previous studies [14, 15, 16, 17]. Notably, two peaks along the long axis of the hexagon are located clearly outside the first Brillouin zone of the surface. The remaining four peaks reside close to the Brillouin zone boundary, but are positioned on its exterior. In addition, we observe Bragg peaks, indicated by yellow circles, which are located at approximately half the wavevectors of previously reported Bragg peaks. We attribute these newly observed Bragg peaks to the primitive CDW wavevectors, determined to be (along the surface unit cell axes defined in Methods) $\mathbf{Q}_{CDW1}=(0.72\pm 0.08,0.02\pm 0.08)$ nm-1 and $\mathbf{Q}_{CDW2}=(0.53\pm 0.08,0.34\pm 0.08)$ nm-1.
The spatially modulated electronic density of states within the CDW state can be expressed as $N_{L}(\mathbf{r})\propto N_{0}sin(\mathbf{Q}_{CDW}\mathbf{r}+\varphi_{CDW})$, with $N_{0}$ the amplitude of the modulations and $\varphi_{CDW}$ the phase. The CDW phase, $\varphi_{CDW}$, is often determined by its interaction with defects. In UTe2, we observe that, generally, maxima in the CDW modulation tend to be pinned at defects (Extended Data Fig. 3). This leads to the formation of domains with CDWs having different phases with respect to the crystal lattice (Extended Data Fig. 4). Furthermore, the phase is independent in different terraces. The CDW modulation locks to the defect distribution on each terrace (Extended Data Fig. 5), demonstrating that it changes in subsequent layers. This suggests that the CDW has a strongly two-dimensional character and is localized at the surface plane.
Band structure of UTe2.
To elucidate the role of the electronic band structure in the CDW formation, we examine the Fermi surface derived from density functional theory (DFT) incorporating Coulomb repulsion $U$ (Fig. 3a, for $U=$2 eV). The Fermi surface is predominantly composed of two nearly two-dimensional tubular bands situated within the $k_{x}$-$k_{y}$ plane, exhibiting a slight warping along the $k_{z}$-axis. These bands display electron-like character along the $k_{x}$ axis and hole-like character along the $k_{y}$ axis bands [34, 33]. The U-5f orbital weight dominates over the whole Fermi surface, with little directional dependence (Fig. 3b). These calculations reproduce well the Fermi surface geometry, although they fail to account for the exceptionally large effective masses observed in UTe2 via de Haas van Alphen and specific heat measurements [35, 36, 37], consistent with observations in other heavy fermion metals [38].
Despite the limitations in reproducing effective masses, the shape of the Fermi surface exhibits a distinctly low-dimensional character, prompting an analysis of the electronic properties expected for the (011) surface. These surface properties are determined by the projection of all isoenergy planes parallel to the surface, i.e. planes perpendicular to the vector normal to the surface plane, ${\bf k}_{\perp}$, ranging from ${\bf k}_{\perp}=0$ to ${\bf k}_{\perp}=\Gamma-S$, where $\Gamma-S$ is the vector joining the center $\Gamma$ and the S point of the bulk Brillouin zone (Fig. 1d,e). This is due to the fact that ${\bf k}_{\perp}$ is not a good quantum number at the surface. The intersections of the bulk Brillouin zone with planes at $k_{\perp}=0$ and $k_{\perp}=|\Gamma-S|$ are indicated by grey rectangles in Fig. 3a. The corresponding Fermi surface contours for the $k_{\perp}=|\Gamma-S|$ plane are shown in Fig. 3d. To identify the $k_{\perp}$ values which contribute most significantly to the tunneling conductance, we calculated the normalized two-dimensional density of states in each plane, using the formula $g_{Norm,\nu}({\bf k}_{\perp})=\frac{1}{L}\int_{L}\frac{dL}{|{\bf\nabla_{k}}E({\bf{k}})|}$, where $L$ is the Fermi surface contour length, $E({\bf{k}})$ the full three-dimensional energy dispersion and $\nu$ the band index (see Extended Data Fig. 6 and Supplementary Information, Section 2 for more details). The results for $g_{Norm,\nu}(k_{\perp})$ for the electron and hole-like bands are shown in Fig. 3c. The largest contributions originate from the planes at the Brillouin zone center ($k_{\perp}=0$) for the hole-like band and boundary ($k_{\perp}=|\Gamma-S|$) for the electron-like band. At $k_{\perp}=|\Gamma-S|$, the Fermi surface contour comprises two parallel lines from the electron-like band. The arrow in Fig. 3d represents the $\mathbf{Q}_{CDW1}$ wavevector observed in our experiment. The near coincidence of this wavevector to the parallel lines suggests that the CDW wavevector along this direction connects near-parallel features in the electron-like band. To dicuss features associated to the $\mathbf{Q}_{CDW2}$ wavevector, we must consider the Fermi contour at $k_{\perp}=0$, which primarily intersects the hole-like band.
The Fermi surface contour of the hole-like band within the $k_{\perp}=0$ plane is shown in Fig. 4 and exhibits significantly greater complexity than the two parallel lines depicted in Fig. 3d, extending well beyond the boundaries of the surface Brillouin zone (indicated in red in Fig. 3d). To represent the two-dimensional Fermi surface contour, we performed a folding operation, mapping the the $k_{\perp}=0$ hole-like band contour into the first surface Brillouin zone by vector addition of reciprocal lattice wavevectors of the surface lattice, denoted as $\overline{\bf{G}}_{S}$, to contour segments located outside the first Brillouin zone. As shown in Fig. 4 we find two vectors connecting band structure segments, oriented at an angle relative to the $k_{x}$-axis. We emphasize that these vectors arise only at the surface after downfolding the bandstructure. These vectors (indicated by light green arrows in Fig. 4) are equivalent by symmetry and coincide with the wavevector $\mathbf{Q}_{CDW2}$ observed in our experiments.
These Fermi surface nesting features are distinct from those associated with enhanced antiferromagnetic fluctuations observed in neutron scattering experiments. The latter occur at a wavevector (0,$\pm$0.93,0) nm-1[28, 27, 29]. But this wavevector has no relation to $\mathbf{Q}_{CDW1}$ or $\mathbf{Q}_{CDW2}$, even if we consider just its projecton to the surface plane (which gives $(\overline{k_{x}}$, $\overline{k_{y}})=(0,0.37)\ \textrm{nm}^{-1}$). Thus, the nesting feature at $\mathbf{Q}_{CDW1}$ does not have a significant influence in the known bulk properties, and the one at $\mathbf{Q}_{CDW2}$ is only found at the surface.
Bulk vs surface electronic properties of UTe2.
While both $\mathbf{Q}_{CDW1}$ and $\mathbf{Q}_{CDW2}$ are discernible on the Fermi surface contours derived from DFT calculations, it is evident that they do not correlate with strong nesting features within the bulk. Determining the actual driving mechanism relating the Fermi surface nesting features identified here and the CDW remains open. A potential origin for the observed modifications in electronic properties at the surface is the U-5f character. The valence state of rare earth and actinide elements is known to vary at surfaces, typically shifting towards a lower f-electron count. [39] Experimental determinations of the U-5f valence in UTe2 using photoemission indicate a tendency towards 5f occupancy near two (U4+) for surface-sensitive and near three (U3+) for bulk-sensitive measurements [30, 40, 41, 42]. In numerous rare earth compounds, CDW formation is found to compete with Kondo screening [43, 44, 45]. CDWs are relatively unfrequent in heavy fermion systems. CDWs consisting of sites with alternating conduction and f-electron character have been proposed to occur within the periodic Anderson model [46]. Our experiment suggests that the reduction of the f-electron valence at the surface and the concomitant decrease in the Kondo temperature might favor the establishment of a CDW.
Note that the present bulk band structure calculations take into account a Coulomb repulsion, but not the Kondo correlations inducing the heavy masses. We observe that a decrease in valence towards U4+ could be associated to enhanced localization and, consequently, enhanced $U$. However, the calculations reveal substantial band structure modifications for decreasing $U$, while exhibiting minimal changes for $U$ values above 2 eV [33]. Thus, the dominance of f-electron character in the nesting features suggests that the CDW could result from f-electron enhanced density of states.
Conclusion.
Leveraging enhanced resolution at low wavevectors in high-resolution STM experiments, we have determined the primitive wavevectors of the CDW in UTe2 and demonstrated that they are located well within the surface Brillouin zone. Furthermore, we established clearly that CDW occurs only at the surface. By comparing our experimental results with band structure calculations that accurately reproduce the Fermi surface geometry, we have identified two potential nesting wavevectors. The valence reduction observed in other surface-sensitive experiments, suggests that surface induced modifications in the configuration of Uranium f-electrons allows for a further enhancement of the electronic susceptibility and the establishment of the CDW at these nesting features.
Figure 1:  | Crystal structure and CDW of UTe2. a Atomic lattice of UTe2. Violet, red and green spheres depict Te(1), Te(2) and U atoms, respectively. The cleaved (011) plane is given in grey. b Atomic resolution topographic image of the UTe2 surface. The atomic lattice positions are marked by dots, with colors corresponding to those in a. The white scale bar represents 1 nm. Te(2) atoms form chains along the crystallographic a-direction. c Schematic phase diagram of UTe2, adapted from Refs. [1, 3]. Phases are labelled as field polarized (FP), superconducting in the FP phase (SCFP) and superconducting in the paramagnetic phase (SCPM). The black dashed line indicates the direction of the applied magnetic field, perpendicular to the cleaved (011) surface. d-f Black lines schematically represent the bulk Brillouin zone. Reciprocal space axes are shown in the bottom left corner of each panel. Blue dots denote selected high symmetry points of the bulk Brillouin zone. The complete set of high-symmetry points for both the bulk and the surface Brillouin zones is presented in Extended Data Fig. 1 and discussed in Supplementary Information, Section 1. Orange dots provide the positions of the Bragg peaks previously reported in Refs. [14, 15, 16, 17]. Yellow dots denote the position of the Bragg peaks of the CDW identified in this study.
Figure 2:  | Tunneling conductance maps of the CDW. Tunneling conductance maps of UTe2 at 10 T a, 15 T b and 20 T c, acquired from three distinct fields of view, and their corresponding Fourier transforms d,e,f. The white scale bars are 5 nm long. The bias voltage for all three maps is close to zero. For tunneling conductance maps as a function of the bias voltage, refer to Extended Data Fig. 2 and Supplementary videos. The white scale bar in the Fourier transforms d-f represents 1 nm-1. The color scale for each map is indicated by the bars on the left. Bragg peaks associated with the atomic lattice are marked by white circles in d,e,f. The surface Brillouin zone is delineated by red lines in d,e,f (see also Extended Data Fig. 1 and Supplementary Infomation, Section 1). Bragg peaks corresponding to charge modulations reported in previous experiments are located slightly outside the surface Brillouin zone boundary and are marked by orange circles. The newly identified CDW wavevectors are indicated by yellow circles in d,e,f.
Figure 3:  | Bulk Fermi surface and Fermi contours at the surface of UTe2. a The Fermi surface is represented as a colored contour, with blue indicating electron-like bands and yellow indicating hole-like bands. The bulk Brillouin zone is delineated by a black line. Reciprocal space planes parallel to the surface, specifically at $k_{\perp}=0$ and $k_{\perp}=|\Gamma-S|$, are shown in grey. b The U 5f character of the bands crossing the Fermi surface is represented using the color scale shown at the top of the figure. c Normalized density of states of the Fermi surface contour $g_{Norm,\nu}({\bf k}_{\perp})=\frac{1}{L}\int_{L}\frac{dL}{|\nabla_{\bf{k}}E|}$, where $L$ is the length of the Fermi surface contour in reciprocal space and $\nu$ is the band index. Blue and yellow colors correspond to the respective band, as in a. d Fermi surface contour at the reciprocal space plane $k_{\perp}=|\Gamma-S|$. The ${\bf k}_{\parallel}$ axis is perpendicular to ${\bf k}_{x}$ within the (011) surface plane in reciprocal space (depicted in grey in a). The surface Brillouin zone is delineated by a red line (see also Extended Data Fig. 1). The wavevector $\mathbf{Q}_{CDW1}$, indicated by a green line, is very close to the nesting wavevector of the electron-like band.
Figure 4:  | Fermi contour of the hole-like band at the surface of UTe2. We depict the atomic Bragg lattice as black points within the reciprocal space surface plane ${\bf k}_{\perp}=0$. The $\overline{{\bf k}_{x}}$ axis corresponds to the crystallographic a-axis, and $\overline{{\bf k}_{y}}$ is perpendicular to $\overline{{\bf k}_{x}}$ and defines the vertical axis of the grey plane in Fig. 3a (see also Methods). The first Brillouin zone of the surface is delineated by red lines and is repeated into adjacent reciprocal space regions. The Fermi contour of the hole band is represented by yellow lines. Black lines demarcate the boundaries of higher-order Brillouin zones. The Brillouin zones are numbered from $1-5$ and are distinguished by different colors. The Fermi contours in higher-order Brillouin zones are folded into the first Brillouin zone by vector addition of a reciprocal lattice wavevector of the surface $\overline{{\bf G_{S}}}$. Light green lines indicate the experimentally determined wavevector $\mathbf{Q}_{CDW2}$.
Methods
Scanning Tunneling Microscopy.
[…]
To determine the three dimensional coordinates of the CDW wavevectors, we utilize the unit vector normal to the (011) surface, $\hat{\bf{k}}_{\perp}=\frac{\bf{k}_{\perp}}{k_{\perp}}=\frac{(0,c,b)}{\sqrt{b^{2}+c^{2}}}$ and two orthogonal unit vectors within the surface plane, $\hat{\bf{k}}_{\perp}$, $\hat{\bf{k}}_{x}=(1,0,0)$ and $\hat{\bf{k}}_{||}=\hat{\bf{k}}_{\perp}\times\hat{\bf{k}}_{x}$. The vectors $\hat{\bf{k}}_{x}$ and $\hat{\bf{k}}_{||}$ define the in-plane coordinate system of the (011) surface. We calculate the in-plane components of the positions of the CDW within the three dimensional bulk Brillouin zone by projecting the vectors providing the positions of the CDW peaks on the surface plane, using ($\overline{k_{x}}$, $\overline{k_{y}}$) =$(\hat{\bf{k}}_{x}\cdot\bf{Q}_{i},\hat{\bf{k}}_{||}\cdot\ \bf{Q}_{i})$, where $\bf{Q}_{i}$ represent the positions of the CDW peaks in the bulk three-dimensional Brillouin zone [23].
The additional CDW Bragg peaks we observe here are also discernible in previous work, for example in Ref. [14] but were likely overlooked. This oversight may have resulted from the obscuring influence of density of states fluctuations at small wavevectors, potentially induced by defects leading to a decreased resolution for small wavevectors. As detailed in the main text, the Bragg peaks observed in this study are located entirely within the first Brillouin zone of the bulk and surface lattices. In contrast, previously observed Bragg peaks are situated well outside the bulk Brillouin zone and marginally outside the surface Brillouin zone.
Band structure calculations.
We performed fully relativistic all-electron density functional theory (DFT) calculations for high-quality UTe2 [55, 56] using the full potential local orbital (FPLO) basis [57] and the generalized gradient approximation (GGA) exchange-correlation functional [58]. To account for the electronic correlations we employed GGA + $U$ calculations with the around mean-field double-counting correction scheme [59]. A $12\times 12\times 12$ $k$-point mesh was utilized. In this study, we select Slater parameters for U $5f$ orbitals as $(F_{0},F_{2},F_{4},F_{6})=(2\mathrm{\;eV},0,0,0)$, corresponding to $\tilde{U}=U-J=2\mathrm{\;eV}$. With these parameters, we obtain a cylindrical Fermi surface characterized by significant U-$5f$ orbital contribution on the Fermi level, consistent with previous DFT studies and with the geometry of the Fermi surface obtained from de Haas van Alphen and angular resolved photoemission measurements measurements [34, 60, 61, 35, 62, 36, 63, 64]. To extract the electronic properties on the surface plane we imported the FPLO output files into Matlab and defined the surface plane and associated cuts using code available in Ref. [49].
Acknowledgments.
We acknowledge discussions with J.P. Sanchez and with G. Lander. Support by the Spanish Research State Agency (PID2023-150148OB-I00, TED2021-130546BI00, PDC2021-121086-I00 and CEX2023-001316-M), the European Research Council PNICTEYES through grant agreement 679080 and VectorFieldImaging Grant Agreement 101069239, the EU through grant agreement No. 871106, by the Comunidad de Madrid through projects projects TEC-2024/TEC-380 “Mag4TIC” and PhD thesis support PIPF-2023/TEC-30683 is acknowledged. We have benefitted from collaborations through EU program Cost CA21144 (superqumap), and from SEGAINVEX at UAM in the design and construction of STM and cryogenic equipment.
Supplementary Information
1 Surface and bulk Brillouin zones
It is useful to remind that the surface Brillouin zone is the Wigner Seitz cell formed by the projection of the bulk Brillouin zone’s $\Gamma$ points onto the surface plane in reciprocal space. In general, the surface Brillouin zone differs from any cross-section of the bulk Brillouin zone defined by the surface plane, i.e. from the bulk Brillouin zone planes at any $k_{\perp}$[65]. For example, in a face-centered-cubic (FCC) crystal, the surface Brillouin zone for the (100) surface is a square. The cut of the bulk Brillouin zone for the plane perpendicular to the direction $\Gamma-X$ at $k_{\perp}=\Gamma-X$ is also a square, but with smaller dimensions than the surface Brillouin zone. At $k_{\perp}=0$, the cut of the bulk Brillouin zone is a parallelogram which is larger than the surface Brillouin zone.
The crystallographic relationship between the bulk and surface Brillouin zones of UTe2 is illustrated in Extended Data Fig. 1. The bulk Brillouin zone is delineated by black lines, with standard high-symmetry point labels. The $\Gamma$ points of adjacent bulk Brillouin zones, defined by the reciprocal lattice vectors ${\bf G}_{S}=\big{(}\pm\frac{1}{a},\pm\frac{1}{b},0\big{)},\ \big{(}\pm\frac{1}{a},0,\pm\frac{1}{c}\big{)},\ \big{(}0,\pm\frac{1}{b},\pm\frac{1}{c}\big{)}\ \mathrm{and}\ \big{(}0,0,\pm\frac{1}{c}\big{)}$, are indicated by blue stars. In Extended Data Fig. 1b, the bulk Brillouin zone is shown projected onto the (011) plane. In this projection, the bulk Brillouin zone points $\Gamma$ and $S$ coincide. The surface Brillouin zone, defined as the Wigner-Seitz cell of the lattice formed by the projected bulk $\Gamma$ points (blue stars) onto the (011) plane, is represented by the red parallelogram in Extended Data Fig. 1b. The high-symmetry points of the surface Brillouin zone are labelled using standard notation for a centered orthorhomic two-dimensional Brillouin zone, namely $\overline{\Gamma}$, $\overline{X}$, $\overline{Y}$ and $\overline{Y^{\prime}}$.
It is important to note that the symmetry of the surface Brillouin zone differs from that of the surface projection of the bulk Brillouin zone. Specifically, points $L_{1}$ and $L_{2}$ are not symmetry-equivalent in the bulk, as shown in Extended Data Fig. 1a. However, the corresponding points in the surface Brillouin zone, $\overline{Y^{\prime}}$, which lie near the projection of $L_{1}$ and $L_{2}$ to the (011) surface, are symmetry-equivalent, as shown in Extended Data Fig. 1b. The two-dimensional surface point group is $D_{1}$. The two-dimensional reciprocal lattice vectors of the surface Brillouin zone $\overline{{\bf G}_{S}}$ join the projections of the bulk $\Gamma$ points to the surface plane. For a comprehensive list of high symmetry point labels refer to Ref. [66].
2 Fermi contours along k
The surface density of states is calculated as the summation of contributions from Fermi contours within each $k_{\perp}$ plane. The tunneling conductance measured by STM is proportional to the surface density of states. However, determining the specific contributions to the tunneling conductance from individual portions of the band structure is challenging, as it depends on the tunneling matrix elements that govern tip-sample coupling. Nevertheless, we can reasonably assume that the $k_{\perp}$ planes exhibiting significant contributions to the density of states also contribute substantially to the tunneling conductance. To quantify these contributions, we calculated the density of states for each $k_{\perp}$ plane and each band, using the expression $g_{\nu,DOS}(k_{\perp})=\int_{L}\frac{dL}{|\nabla_{\bf{k}}E|}$, with the results presented in Extended Data Fig. 6a-c. We also calculated the density of states normalized by the respective contour length, $g_{Norm,\nu}(k_{\perp})=\frac{1}{L}\int_{L}\frac{dL}{|\nabla_{\bf{k}}E|}$, as shown in Extended Data Fig. 6d-f. The normalized density of states reflects the contribution from band dispersion, independent of contour variations with changing $k_{\perp}$. Our analysis indicates that the largest contributions to the tunneling conductance arise from $k_{\perp}\approx 0$ for the hole-like band and $k_{\perp}\approx|\Gamma-S|$ for the electron-like band.
3 Origin of the CDW in $\alpha-$U, in dichalcogenides and in heavy fermions
3.1 Electron-phonon driven CDW.
Most models for CDWs consider the cooperative interplay of lattice and electronic interactions [67, 68, 69, 70, 71]. The basic mechanism is the Peierls instability of a one-dimensional lattice and is characterized by perfect nesting of Fermi points at opposite wavevectors $k_{F}$ and $-k_{F}$. This results in a dimerized lattice and an insulating ground state at low temperatures. However, perfect Fermi surface nesting is rarely observed, even in quasi one-dimensional systems [72]. Many CDW materials remain metallic in the ground state. A notable example is $\alpha-$U, which displays a metallic CDW accompanied by a structural transition [73, 74, 75].
Such electron-phonon driven CDWs can be viewed as a consequence of the dynamical susceptibility $\chi(\bm{q})$ of the electronic system presenting a peak in its real part $\chi^{\prime}(\bm{q})=Re(\chi(\bm{q}))$ at $\bm{q}=\mathbf{Q}_{CDW}$. This favors charge modulations at $\mathbf{Q}_{CDW}$. Nesting gives a peak on the imaginary part of the susceptibility at wave vectors joining nested Fermi surface portions at $\bm{q}=2\mathbf{k}_{F}$, $\chi^{\prime\prime}(\bm{q})=Im(\chi(\bm{q}))$, which is not necessarily sufficiently strong to be reflected in $Re(\chi(2\mathbf{k}_{F}))$ and lead to a CDW [72, 76].
The real component of the electronic susceptibility $\chi^{\prime}(\bm{q})$ provides the energy change induced by a lattice perturbation with wavevector $\bm{q}$, often compensated by a change in the atomic positions [77, 78]. We can write
$Re\left(\chi(\bm{q})\right)$ $=\chi^{\prime}(\bm{q})=\sum_{\bm{k}}\frac{f(\epsilon_{\bm{k}})-f(\epsilon_{\bm{k+q}})}{\epsilon_{\bm{k}}-\epsilon_{\bm{k+q}}}$ (1)
$Im\left(\chi(\bm{q})\right)$ $=\chi^{\prime\prime}(\bm{q})\quad\text{and}\quad\lim_{\omega\rightarrow 0}\frac{\chi^{\prime\prime}(\bm{q},\omega)}{\omega}=\sum_{\bm{k}}\delta(\epsilon_{\bm{k}}-\epsilon_{F})\delta(\epsilon_{\bm{k+q}}-\epsilon_{F})$
where $f(\epsilon_{\bm{k}})$ is the Fermi-Dirac occupation function. The change in electronic energy depends only on the real part of $\chi(\bm{q})$, $\chi^{\prime}(\bm{q})$.
The imaginary part of the susceptibility, or, more precisely, $\lim_{\omega\rightarrow 0}\frac{\chi^{\prime\prime}(\bm{q})}{\omega}$ is also called the nesting function. It can be readily seen that it selects only energies at the Fermi level, and presents a peak at a given $\bm{q}$ when many wave-vectors ${\bm{k}}$ on the Fermi surface have the same Fermi energy as ${\bm{k+\bm{q}}}$, hence when $\bm{q}$ is a Fermi surface nesting vector. The nesting wavevector and the wavevector of an eventual CDW are however not directly linked. Only a peak in $\chi^{\prime}(\bm{q})$ allows for the absence of screening at a given wavevector required for the establishment of a CDW.
We need to realize that $\chi^{\prime}(\bm{q})$ behaves differently than $\chi^{\prime\prime}(\bm{q})$ because $\chi^{\prime}(\bm{q})$ depends on energy levels below and above the Fermi level, whereas $\chi^{\prime\prime}(\bm{q})$ is just a Fermi surface property. Hence, $\chi^{\prime}(\bm{q})$ can have a peak without nesting and, conversely, nesting does not imply a peak in $\chi^{\prime}(\bm{q})$.
It is interesting to note that a peak in $\chi^{\prime}(\bm{q})$ is found in a simple one-dimensional model with one completely filled band [72]. We can consider a one-dimensional insulator, with a filled band of dispersion $\epsilon_{1}(\bm{k})=cos(ka)$, and an empty valence band of dispersion $\epsilon_{2}(\bm{k})=\epsilon+cos(ka)$, with $\epsilon>2$. Then, $\chi^{\prime}(\bm{q})$ displays a peak at $q=\frac{\pi}{a}$. Only terms joining different bands contribute to Eq.1, as these are the only ones with non-zero numerator $f(\epsilon_{1}(\bm{k}))-f(\epsilon_{2}(\bm{k+q}))=1$. For these states, there is always a finite energy difference $\Delta E=\epsilon_{1}(\bm{k})-\epsilon_{2}(\bm{k+q})$. We see that if the energy difference becomes small, $\frac{f(\epsilon_{\bm{k}})-f(\epsilon_{\bm{k+q}})}{\epsilon_{\bm{k}}-\epsilon_{\bm{k+q}}}$ tends towards $f^{\prime}\left(1/2(\epsilon_{\bm{k}}+\epsilon_{\bm{k+q}})\right)$. The derivative of $f(\epsilon)$ is peaked at $\epsilon_{F}$, hence the biggest contribution is for $\epsilon_{\bm{k}}$ and $\epsilon_{\bm{k+q}}$ on different sides of $\epsilon_{F}$, and equally away from the Fermi level. For a larger energy difference, the ratio $\frac{f(\epsilon_{\bm{k}})-f(\epsilon_{\bm{k+q}})}{\epsilon_{\bm{k}}-\epsilon_{\bm{k+q}}}$ remains maximum when $f(\epsilon_{\bm{k}})-f(\epsilon_{\bm{k+q}})$ is largest, hence still when $(\epsilon_{\bm{k}},\epsilon_{\bm{k+q}})$ remains centered around $\epsilon_{F}$. The condition $(\epsilon_{\bm{k}},\epsilon_{\bm{k+q}})$ centered around $\epsilon_{F}$ means that both are equidistant from $\epsilon_{F}$ but on opposite sides (one above, the other below). Therefore, close to the Fermi level, we can have states for which $\epsilon_{\bm{k}}=\epsilon_{\bm{k+q}}$ (maximum of $f^{\prime}$, the derivative of $f$) and $\frac{d\epsilon}{dk}(\bm{k})=v_{F}(\bm{k})$ is exactly opposite to $\frac{d\epsilon}{dk}(\bm{k+q})=v_{F}(\bm{k+q})$. When such two levels remain centered around $\epsilon_{F}$, the contribution to $\chi^{\prime}(\bm{q})$ is maximized [79, 80, 72]. Naturally, nested regions give some contribution to $\chi^{\prime}(\bm{q})$, but a peak develops most strongly if the condition $v_{F}(\bm{k})=-v_{F}(\bm{k+q})$ is satisfied. Hence, for the simple one-dimensional model considered, there is a peak in $\chi^{\prime}(\bm{q})$ at $q=\frac{\pi}{a}$. The conditions leading to a peak in $\chi^{\prime}(\bm{q})$, in particular $v_{F}(\bm{k})=-v_{F}(\bm{k+q})$, remain for more realistic band structures [79, 80, 72].
Therefore, in $\alpha-$U, the CDW can be explained by considering states of equal energy above and below $E_{F}$, for which $v_{F}(\bm{k})=-v_{F}(\bm{k+q})$. Indeed, nesting of narrow f-electron U derived bands were first identified as a possible source of the CDW [81]. However, the observed pressure dependence of the CDW was found to be incompatible with the near absence of pressure induced effects on the Fermi surface nesting properties [82]. It was then shown that phonon softening, observed in neutron scattering and only obtained in calculations after incorporating fully itinerant U-5f electrons, is instead responsible for the CDW [82]. Subsequently, it was shown that nesting features of the Fermi surface do not explain the CDW [76]. Instead, portions of the band structure not lying exactly at the Fermi level and for which $v_{F}(\bm{k})=-v_{F}(\bm{k+q})$ produce the relevant electronic contributions to the peak in $\chi^{\prime}(\bm{q})$ which leads to the CDW in $\alpha-$U [79, 76].
3.2 CDWs in heavy fermions and in UTe2.
The decrease in the Kondo temperature suggested by our measurements at the surface reduces the relevance of Kondo screening and could favor the establishment of a peak in $\chi^{\prime}(\bm{q})$. The nesting feature we observe at $\mathbf{Q}_{CDW1}$ connects the same electron-like band and thus naturally allows for $v_{F}(\bm{k})=-v_{F}(\bm{k+q})$. However, calculations and experiments show that the largest peak in the bulk susceptibility is rather located along another direction, which coincides with the wavevector where antiferromagnetic fluctuations are observed in neutron scattering[28, 27, 29].
Furthermore, the same electron-like band has a strong U-5f character close to the Fermi level, which decreases for lower energies towards the bottom of the band. At the band bottom, approximately at $E=-0.4$ eV on the projected electron-like band at the surface (leading to $\mathbf{Q}_{CDW1}$ at the Fermi level), the U-6d character is much more pronounced and provides a strong peak in the density of states reminiscent of the peak found for a one-dimensional electronic band dispersion [42].
It is interesting to note that the appearance of the CDW in our STM experiments is highly dependent on the bias voltage. As we show in the Extended Data Figure 2 and in the supplementary videos, the peaks at $\mathbf{Q}_{CDW1}$ and $\mathbf{Q}_{CDW2}$ change their intensity with energy. Generally, there is a tendency to a stronger intensity for negative energies (below $E_{F}$).
At the same time, impurities are also much better visible at negative energies. The local density of states increases at impurities and at the crests of the CDW modulations, as can be readily seen in Extended Data Fig. 3. Given that these changes are mostly observed at negative energies, is tempting to associate them to variations in the U-6d character of the local density of states. To address this issue, we can calculate the cross-correlation of tunneling conductance maps, defined as [83]:
$CC=\frac{\sum_{x,y}[G(V_{1},x,y)-\overline{G(V_{1},x,y)}][G(V_{2},x,y)-\overline{G(V_{2},x,y)}]}{N_{x}N_{y}\sigma_{V_{1}}\sigma_{V_{2}}}$ (2)
at different bias voltages $V_{1}$ and $V_{2}$ ($\sigma_{V_{i}}$ being the standard deviation of all the conductance values found in the map, $N_{x}$, $N_{y}$ the number of points in each direction, $G(V_{i},x,y)$ the tunneling conductance maps and $\overline{G(V_{i},x,y)}$ the average of the tunneling conductance at the bias voltages $V_{i}$).
For patterns in the STM images dominated by a modulation in the atomic positions, we expect to find $CC\approx 1$ for all bias voltages. For patterns in the STM images connected to a Peierls-like CDW with a gap opened in the density of states, the change in $\varphi_{CDW}$ with energy is of $\pi$ accross the gap, due to the opposed character of the band dispersion at both sides of the gap, leading to $CC\approx-1$. We find values around $CC=-0.2$ for bias voltages lying above ($+$5 mV) and below (--10 mV) zero bias, suggesting that the observed features are not atomic displacements, but electronic density modulations. What is more, the patterns strongly change their intensity as a function of the bias voltage, generally leading to $CC<0$ when comparing positive and negative bias voltages. This is difficult to understand within a simple phonon-mediated CDW with a gap opening. Instead, this suggests that spatial variations of the correlated properties of the f-electron band structure lead to the spatial variations in the local density of states.