### v2

# Surface charge density wave in UTe2

Surface charge density wave in UTe2
Pablo García Talavera Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Miguel Águeda Velasco Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Makoto Shimizu Department of Physics, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan. Beilun Wu Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Óscar Bou Marqués Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Georg Knebel University Grenoble Alpes, CEA, IRIG-PHELIQS, F-38000 Grenoble, France. Midori Amano Patino University Grenoble Alpes, CEA, IRIG-PHELIQS, F-38000 Grenoble, France. Univ. Grenoble Alpes, CNRS, Grenoble INP, Institut Néel, 38000 Grenoble, France Univ. Grenoble Alpes, CEA, IRIG, Pheliqs, 38000 Grenoble, France. Gerard Lapertot University Grenoble Alpes, CEA, IRIG-PHELIQS, F-38000 Grenoble, France. Jacques Flouquet Institute for Materials Research, Tohoku University, Ibaraki 311-$13$, Japan. Jean Pascal Brison University Grenoble Alpes, CEA, IRIG-PHELIQS, F-38000 Grenoble, France. Dai Aoki Institute for Materials Research, Tohoku University, Ibaraki 311-$13$, Japan. Youichi Yanase Department of Physics, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan. Edwin Herrera Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Isabel Guillamón Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Hermann Suderow Laboratorio de Bajas Temperaturas y Altos Campos Magnéticos, Unidad Asociada UAM-CSIC, Departamento de Física de la Materia Condensada and Instituto Nicolás Cabrera, Universidad Autónoma de Madrid, E-28049 Madrid, Spain. Condensed Matter Physics Center (IFIMAC), Universidad Autónoma de Madrid, E-28049 Madrid, Spain.
The spatially uniform electronic density characteristic of a metal can become unstable at low temperatures, leading to the formation of charge density waves (CDWs). These CDWs, observed in dichalcogenides, cuprates, and pnictides arise from features in the atomic lattice and its interaction with the electronic band structure that facilitate charge ordering. However, CDWs are rarely observed in presence of Kondo screening and heavy fermion quasiparticles. The heavy fermion topological superconductor candidate UTe2 presents a notable exception, exhibiting a CDW whose origin remains elusive. Here we report high resolution Scanning Tunneling Microscopy (STM) experiments that reveal the primitive wavevectors of the CDW in UTe2. This allows for a refined identification of the nesting wavevectors in the electronic bandstructure. Although these wavevectors have no specific influence on the bulk properties, for example on antiferromagnetic fluctuations, they cause the interactions leading to the CDW at the surface. The heavy fermion hybridization pattern is spatially modulated specifically at the nesting wavevectors, suggesting that surface induced modifications in the U 5f electron valence enable a novel form of purely electron-driven charge ordering.
The CDW in UTe2.
The unconventional superconductivity observed in UTe2 exhibits characteristics strongly suggestive of spin-triplet Cooper pairing. The superconducting phase diagram, featuring an exceptionally high critical field and reentrant behavior at extreme magnetic fields[1, 2, 3], anomalous field dependence of the Knight-shift[4, 5] and multiple transitions between superconducting phases[6, 7, 8, 9, 10, 11, 12, 4], poses significant challenges to explanations based on conventional s-wave BCS theory. A relevant question concerns the nature of the normal state parent to the remarkable superconducting properties of UTe2.
Scanning Tunneling Microscopy (STM) studies have revealed the presence of a charge density wave (CDW)[13, 14, 15, 16, 17]. Analogous to the cuprates[18, 19, 20], the CDW is linked to a spatially modulated superconducting phase, termed pair density wave (PDW)[14, 21]. However, the underlying mechanism linking the CDW to electronic properties remains poorly understood. Notably, structural studies have shown the absence of detectable lattice distortions characteristic of conventional CDW formation[22, 23], and macroscopic measurements, such as specific heat, ultrasound attenuation or thermal expansion have failed to identify signatures of a phase transition associated with the CDW[24, 25]. Moreover, the hitherto observed CDW wavevectors lie outside the first Brillouin zone and have no relationship with features of the bulk band structure[13, 14, 15, 16, 26, 27, 28, 29, 30, 31, 32]. Here, we report new high resolution millikelvin STM measurements focusing at high magnetic fields to address as closely as possible the behavior of the normal phase. We demonstrate that the previously identified wavevectors are not primitive wavevectors of the CDW lattice and unveil the electronic origin of the CDW.
Primitive wavevectors of the CDW of UTe2.
The cleaved surface of UTe2, depicted in grey in Fig. 1a, corresponds to the (011) plane of the orthorhombic crystal structure (space group #71, $Immm$). The STM topography image in Fig. 1b reveals chains formed by Te(2) atoms, with chains of Te(1) located slightly below. Uranium (U) chains are located between the Te(1) and Te(2) surface chains. The atomic arrangement at the surface defines the surface centered rectangular unit cell, indicated by the solid blue lines in Fig. 1b. The primitive crystal surface unit cell is represented by dashed blue lines in Fig. 1b. We applied a magnetic field perpendicular to the surface, along a direction within the superconducting phase diagram at $23.7^{\circ}$ from the b axis towards the c axis direction, as shown in Fig. 1c. This magnetic field direction aligns with the onset of the very high field superconducting phase above 35 T[1, 2, 3]. The bulk Brillouin zone, with standard high symmetry point labels, is shown in black lines in Fig. 1d-f. The previously reported CDW Bragg peaks are shown in orange, and lie outside of both the bulk and the surface first Brillouin zones (more details in Extended Data Fig. 1 and in Supplementary Information, Section 1). In contrast, the Bragg peaks observed in this study, depicted in yellow, are located within the first Brillouin zones of both the bulk and the surface lattices.
We present atomic resolution tunneling conductance maps acquired near zero bias from 10 T to 20 T in Fig. 2a-c, illustrating data from three different fields of view. The corresponding Fourier transforms are shown in Fig. 2d-f. Tunneling conductance maps at different bias voltages are provided in Extended Data Fig. 2 and in Supplementary videos. We identify Bragg peaks associated with the atomic lattice, marked with white circles, and an additional set corresponding to the CDW. The surface Brillouin zone is delineated by a red hexagon in Fig. 2d-f. The CDW peaks indicated by orange circles are consistent with those reported in previous studies[13, 14, 15, 16]. Notably, two peaks along the long axis of the hexagon are located clearly outside the first Brillouin zone of the surface. The remaining four peaks reside close but outside the Brillouin zone boundary. In addition, we observe Bragg peaks, indicated by yellow circles, which are located at approximately half the wavevectors of previously reported Bragg peaks. We attribute these newly observed Bragg peaks to the primitive CDW wavevectors, determined to be (along the surface unit cell axes defined in Methods) $\bm{Q}_{CDW1}=(0.72\pm 0.08,0.02\pm 0.08)$ nm-1 and $\bm{Q}_{CDW2}=(0.53\pm 0.08,0.34\pm 0.08)$ nm-1.
The electronic density of states spatially modulated by the CDW can be expressed as $N_{L}(\bm{r})\propto N_{0}\sin(\bm{Q}_{CDW}\cdot\bm{r}+\varphi_{CDW})$, with $N_{0}$ the amplitude of the modulations and $\varphi_{CDW}$ the phase. We find that the phase $\varphi_{CDW}$ is independent in different terraces (Supplementary Information Section 2 and Extended Data Figs. 3,4,5). This suggests that the CDW has a strongly two-dimensional character and is localized at the surface plane.
Band structure of UTe2.
To elucidate the role of the electronic band structure in the CDW formation, we examine the Fermi surface derived from density functional theory (DFT) incorporating Coulomb repulsion $U$ (Fig. 3a, for $U=$ 2 eV). The Fermi surface is predominantly composed of two nearly two-dimensional tubular bands situated within the $k_{x}$-$k_{y}$ plane, exhibiting a slight warping along the $k_{z}$-axis. These bands display electron-like character along the $k_{x}$ axis and hole-like character along the $k_{y}$ axis[33, 32]. The U 5f orbital weight dominates over the whole Fermi surface, with little directional dependence (Extended Data Fig. 7b). These calculations reproduce well the Fermi surface geometry, although they fail to account for the exceptionally large effective masses observed in UTe2 via de Haas van Alphen and specific heat measurements[34, 35, 36, 37], consistent with results in other heavy fermion metals[38].
Despite the limitations in reproducing effective masses, it is useful to analyze the electronic properties expected for the (011) surface. These surface properties are determined by the projection of all isoenergy planes parallel to the surface, i.e. planes perpendicular to $\bm{k}$$_{\perp}$, the vector normal to the surface plane, ranging from ${\bm{k}}_{\perp}=0$ to ${\bm{k}}_{\perp}=\Gamma-S$, where $\Gamma-S$ is the vector joining the center $\Gamma$ and the $S$ point of the bulk Brillouin zone (Fig. 1d,e). This is due to the fact that $\bf k$$_{\perp}$ is not a good quantum number at the surface. The intersections of the bulk Brillouin zone with planes at $k_{\perp}=0$ and $k_{\perp}=|\Gamma-S|$, i.e. through the $\Gamma$ and $S$ points, are shown by grey rectangles in Fig. 3a. The corresponding Fermi surface contours for the $k_{\perp}=|\Gamma-S|$ plane are shown in Fig. 3b. To identify the $k_{\perp}$ values which contribute most significantly to the tunneling conductance, we calculated the normalized two-dimensional density of states in each plane, using the formula $g_{Norm,\nu}(k_{\perp})=\frac{1}{L}\int_{L}\frac{dL}{|\nabla_{\bm{k}}E({\bm{k}})|}$, where $L$ is the Fermi surface contour length, $E(\bm{k})$ the full three-dimensional energy dispersion and $\nu$ the band index (see Extended Data Fig. 6 and Supplementary Information, Section 3 for more details). The results for $g_{Norm,\nu}(k_{\perp})$ for the electron and hole-like bands are shown in Fig. 3c. The largest contributions originate from the planes at the Brillouin zone center ($k_{\perp}=0$) for the hole-like band and boundary ($k_{\perp}=|\Gamma-S|$) for the electron-like band. At $k_{\perp}=|\Gamma-S|$, the Fermi surface contour comprises two parallel lines from the electron-like band. The arrow in Fig. 3b represents the $\bm{Q}_{CDW1}$ wavevector observed in our experiment. The near coincidence of this wavevector to the parallel lines suggests that the CDW wavevector along this direction connects near-parallel features in the electron-like band. To discuss features associated to the $\bm{Q}_{CDW2}$ wavevector, we must consider the Fermi contour at $k_{\perp}=0$, which primarily intersects the hole-like band.
The Fermi surface contour of the hole-like band within the $k_{\perp}=0$ plane is shown in Fig. 3d and exhibits significantly greater complexity than the two parallel lines depicted in Fig. 3b, extending well beyond the boundaries of the surface Brillouin zone (indicated in red in Fig. 3b,d). To represent the two-dimensional Fermi surface contour, we performed a folding operation, mapping the $k_{\perp}=0$ hole-like band contour into the first surface Brillouin zone by vector addition of reciprocal lattice wavevectors of the surface lattice, denoted as $\overline{\bm{G}}_{S}$, to contour segments located outside the first Brillouin zone. As shown in Fig. 3d we find two vectors connecting band structure segments, oriented at an angle relative to the $k_{x}$-axis. We emphasize that these vectors arise only at the surface after downfolding the bandstructure. These vectors (indicated by dark green arrows in Fig. 3d) are equivalent by symmetry and coincide with the wavevector $\bm{Q}_{CDW2}$ observed in our experiments.
These Fermi surface nesting features are distinct from those associated with enhanced antiferromagnetic fluctuations observed in neutron scattering experiments. The latter occur at a wavevector ($\pm0$.93,0) nm-1[27, 26, 28]. But this wavevector has no relation to $\bm{Q}_{CDW1}$ or $\bm{Q}_{CDW2}$, even if we consider just its projection to the surface plane (which gives $(\overline{k_{x}}$, $\overline{k_{y}})=(0,0.37)\ \textrm{nm}^{-1}$). Thus, the nesting feature at $\bm{Q}_{CDW1}$ does not have a significant influence in the known bulk properties, and the one at $\bm{Q}_{CDW2}$ is only found at the surface as it arises from a feature obtained by downfolding the surface Brillouin zone. While both $\bm{Q}_{CDW1}$ and $\bm{Q}_{CDW2}$ are discernible on the Fermi surface contours derived from DFT calculations, they do not correlate with strong nesting features within the bulk.
CDW and the heavy fermion density of states in UTe2.
A phonon-mediated CDW appears in $\alpha$-U (see Supplementary Information, Section 4). However, in rare earth compounds, CDW formation is found to compete with Kondo screening[39, 40, 41]. CDWs are indeed relatively unfrequent in heavy fermion systems. This, together with the absence of indications for a CDW from structural studies in UTe2[22, 23] and the distinct heavy masses not reproduced by DFT[35, 36, 37], prompts us to question the eventual modifications suffered by the electronic band structure at the surface.
The heavy fermion quasiparticles lead to a dip-peak structure in the tunneling conductance of UTe2, which is due to co-tunneling into heavy and light quasiparticles and the hybridization gap opening[42, 43, 44]. In Fig. 4a we show the tunneling conductance on flat terraces in UTe2, together with a fit to the usual expression for heavy fermion low energy hybridization, described in the Methods section. We find that the shape of the dip-peak structure in the tunneling conductance changes together with the CDW (light and dark green crosses on the inset of Fig. 4a), being compatible with local changes in the hybridization strength by about 10% (Extended Data Table 1 and Methods).
[…]
The concept of a spatial variation of the hybridization leading to periodic charge oscillations has been theoretically explored and experimentally observed, with charge modulations at nesting wavevectors that are close to those of the high energy unhybridized band structure. However, these remained local around Kondo holes and impurities[43, 44]. A spatially extended hybridization wave was theoretically proposed within the periodic Anderson model and as a novel form of low temperature electronic ordering in the hidden order state of URu2Si2[49, 50, 51, 52].
Conclusion.
Leveraging enhanced resolution at low wavevectors in high-resolution STM experiments, we have determined the primitive wavevectors of the CDW in UTe2 and demonstrated that they are located well within the surface Brillouin zone. Furthermore, we established clearly that CDW occurs only at the surface. By comparing our experimental results with band structure calculations that accurately reproduce the Fermi surface geometry, we have identified two nesting wavevectors matching the primitive wavevectors of the CDW. Furthermore, the behavior observed at steps suggests that surface-induced modifications in the valence of Uranium f-electrons drives a reduction of the Kondo temperature and allows the establishment of a periodic modulation of the hybridization pattern at the nesting features.
These nesting features do not produce charge or spin ordering in the bulk. However, they are visible in several experiments addressing both surface and bulk properties. For example, very recent experiments have found modulations of the in-gap superconducting density of states at some of the primitive CDW wavevectors [53], adding to the previously observed superconducting PDW associated to the CDW [14]. Furthermore, calculations hint at possible nodes in the superconducting order parameter at the primitive CDW wavevectors [54, 55]. Additionally, the superconducting coherence length anisotropy leads to strongly elongated vortex cores, compatible with the nearly one-dimensional character of the band giving $\bm{Q}_{CDW1}$ [32, 56, 57, 58]. This suggests that the nesting features which lead to the CDW at the surface have also a profound influence on the bulk properties of UTe2.
Acknowledgments.
We acknowledge discussions with J.P. Sanchez and with G. Lander. Support by the Spanish Research State Agency (PID2023-150148OB-I00, TED2021-130546B-I00, PDC2021-121086-I00 and CEX2023-001316-M), the European Research Council PNICTEYES through grant agreement 679080 and VectorFieldImaging Grant Agreement 101069239, the EU through grant agreement No. 871106, by the Comunidad de Madrid through projects projects TEC-2024/TEC-380 “Mag4TIC” and PhD thesis support PIPF-2023/TEC-30683 and PIPF-2023/TEC-30853 are acknowledged. M.A.V., E. H., I.G and H. S. acknowledge the QUASURF project (ref. SI4/PJI/2024-00199) funded by the Comunidad de Madrid through the agreement to promote and encourage research and technology transfer at the Universidad Autónoma de Madrid. We have benefited from collaborations through EU program Cost CA21144 (superqumap), and from SEGAINVEX at UAM in the design and construction of STM and cryogenic equipment.
Author contributions.
P.G.T. performed the experiments with the supervision of I.G. and E.H., and support by M.A.V., P.G.T. and M.A.V. worked together and with M.S. and O.B.M. on data treatment and analysis. M.S. and Y.Y. performed the band structure calculations. M.A.P., G.P. and D.A., M.A.V. and G.L. provided samples and helped to mount these in the STM, together with B.W., G.K., J.F. and J.P.B., G.K., J.F. and J.P.B. and Y.Y. provided guidance about the properties of UTe2 and the interpretation of the results and co-supervised the work, with E.H., I.G. and H.S. who wrote the manuscript, together with P.G.T. All authors contributed to the manuscript.
Data availability.
All data and calculations are available upon request.